2 workers and some handheld computers that you can use to take inventory at a warehouse. There are Imagine you have some computer can together inventory 150 tems per hour Three workers And four or more workers sharing a computer can together inventory fewer than 160 items per hour. Computers cost $100 each and shing returns to taking inventory. If one worker uses one computer, he can 100 items per hour. Two workers sharing a sharing a computer can together inventory 160 items per hour you must pay each worker $25 per hour Instructions: Round your answers to 2 d a If you assign one worker per computer, what is the cost of Defenencesb. b. What f you assign two workers per computer? What is the cost of inventorying a single tem c. What if you assign three? d. How many workers per computer should you assign # you wish to minimize the cost of ーーーコworkets) per computer 。Type here to search
Answer #1

one work with one computer can produce 100 items per hour

labor price of one worker =25/hour

cost of computer= fixed cost=100

total cost=125

total items produced = 100

so price per items=125/100=1.25

for two workers

items producced=150

cost of computer=100

cost of worker=25*2=50

total cost=150

price per item =150/150= $1

C. for 3 workers

items produced=160

cost of computer=100

cost of workers = $25*3=$75

total cost= 175

cost per item= 175/160 = $1.09

As we can see cost per item is least for 2 workers so we would employ two workers